OpenStudy (sleepyjess):

@ankit042 Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

OpenStudy (eric_d):

Find the common diff, d

OpenStudy (ankit042):

for arithmetic seq nth term is a(n)=a(0)+(n-1)d

OpenStudy (ankit042):

a(14) = a(0)+13d a(15)= a(0)+14d subtract the two equation to get d

OpenStudy (sleepyjess):

d = 42

OpenStudy (ankit042):

yes!